OpenStudy (anonymous):

sketch the solid described by the given inequalities (cylindrical coordinates) 0(

OpenStudy (anonymous):

\[\huge 0 \le \theta \le \frac{\pi}{2}\]\[\huge r \le z \le 2\] *idk solution, only type this so it looks clearer*

OpenStudy (ganpat):

i can give a try.. but no assurance..

OpenStudy (ganpat):

 :D